OpenStudy (anonymous):

A 20 loudspeaker and an 12 loudspeaker are connected in parallel across the terminals of an amplifier. Assuming the speakers behave as resistors, calculate the equivalent resistance of the two speakers.

OpenStudy (ipwnbunnies):

For parallel resistance: \[\frac{1}{Rtotal} = \frac{1}{R1} + \frac{1}{R2} +...+\frac{1}{RN}\]